OpenStudy (anonymous):

HELP ALGEBRA PLEASE Write the equation of the line that passes through (–1, 5) and has a slope of 3 in point-slope form.

OpenStudy (anonymous):

@happyrosy

OpenStudy (anonymous):

@triciaal @Zarkon @whpalmer4 @undeadknight26 @jigglypuff314 @Vocaloid

undeadknight26 (undeadknight26):

Oh man I for got point slope for is that y = mx + b?

OpenStudy (anonymous):

x + 1 = 3(y – 5) x – 5 = 3(y + 1) y – 5 = 3(x + 1) y + 1 = 3(x – 5)

undeadknight26 (undeadknight26):

Nerp totally forgot that lol

jigglypuff314 (jigglypuff314):

point slope form would be y - k = m(x - h) where m = slope and point (h, k) is a point on the line

jigglypuff314 (jigglypuff314):

so you can plug in m = 3 h = (-1) k = 5

OpenStudy (anonymous):

so y-5=3(x-1)

OpenStudy (anonymous):

Would it be c?

jigglypuff314 (jigglypuff314):

be careful with plugging in the (-1) for h! and yes, C would be correct :)

OpenStudy (anonymous):

Thanks so much. but how does the negative 1 turn into a positive 1.

jigglypuff314 (jigglypuff314):

because in the formula: \( y - k = m(x - h)\) ^ the minus here turns the (-1) that you plug in into (+1) [negative times negative = positive]

OpenStudy (anonymous):

Thank you!